can you help me solve this problem:
directions are: solve each system by elimination.
-4x-2y=18
-6x-2y=22
jnh_rayven@yahoo.com-a concerned parent

Answer by checkley77(12844)

-4x-2y=18
-6x-2y=22 subtract.
-----------------
2x=-4
x=-4/2
x=-2 ans.
-4*-2-2y=18
8-2y=18
-2y=18-8
-2y=10
y=10/-2
y=-5 ans.
Proof:
(-6*-2)-(2*-5)=22
12+10=22
22=22